# Yoga for Fourier–Mukai partnership

### v2

@@ 4. Fully faithfulness & equivalences
By Proposition˜3.8, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ are adjoint functors on $D_{\operatorname{qc}}$. Choose $E\in D_{\operatorname{qc}}(Y_{1})$. We want to show that $\mathbf{L}t_{1}^{\ast}(\eta_{E})$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for this adjunction.
Let $\beta\colon\mathbf{L}t^{\ast}_{1}E\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ be a morphism in $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for any $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$. Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$, making the following diagram commute,
$\mathbf{L}t^{\ast}_{1}E$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{\beta}$$\scriptstyle{\beta^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
where $\xi^{t_{1}}$ is the counit of derived pullback and pushforward along $t_{1}$. Furthermore, by the universal property of the counit, the morphism $\beta^{\prime\prime}$ is the unique morphism such that $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$. Similarly, using that $\Phi_{K}$ is left adjoint to $\Phi_{K^{\prime}}$ and the natural isomorphism $\alpha$ of Lemma˜4.1, there is a $\beta^{\prime\prime\prime}$ giving a diagram
[…]
which commutes. Pulling back this diagram via $\mathbf{L}t_{1}^{\ast}$, we obtain
$\beta^{\prime}=\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
which tells us $\beta$ factors through $\mathbf{L}t_{1}^{\ast}(\eta_{E})$.
It remains to show that the above factorization for $\beta^{\prime}$ is unique. So, suppose there is a morphism $h\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ such that $\beta=h\circ\mathbf{L}t_{1}^{\ast}(\eta_{E})$. Once again, using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$, we have a morphism
$h^{\prime}\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$
giving a commutative diagram
$\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{h}$$\scriptstyle{h^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
with $\xi^{t_{1}}$ as above and $h^{\prime}=\mathbf{L}t^{\ast}_{1}h^{\prime\prime}$ for a morphism $h^{\prime\prime}\colon\Phi_{K^{\prime}}\Phi_{K}(E)\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$. Here, $h^{\prime\prime}$ is the unique morphism such that $h=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}h^{\prime\prime}$.
Tying things together, we see that
$\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}\circ\eta_{E})=(\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}))\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=h\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=\beta.$
However, by the work above, we know that
$\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\beta^{\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
So, from uniqueness, it follows that
$h^{\prime\prime}\circ\eta_{E}=(\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}\circ\eta_{E}.$
As $\Phi_{K}$ is fully faithful on $D^{b}_{\operatorname{coh}}$, $\eta_{E}$ must be an isomorphism, which implies $h^{\prime\prime}=(\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}$. Consequently, we have
$h=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}h^{\prime\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}),$
and hence, uniqueness as desired to complete the proof. ∎
Lemma 4.3.
Let $F\colon\mathcal{S}_{1}\leftrightarrows\mathcal{S}_{2}\colon G$ be an adjoint pair of exact functors between compactly generated triangulated categories. Assume $G$ preserves small coproducts. Denote by $\eta\colon 1\to GF$ for the unit of the adjunction. Then the collection $\mathcal{T}$ in $\mathcal{S}_{1}$ consisting of objects $E$ for which $\eta_{E}$ is an isomorphism is a triangulated subcategory of $\mathcal{S}_{1}$ which is closed under small coproducts. Moreover, $F$ is fully faithful if, and only if, $\eta_{E}$ is an isomorphism for every $E\in\mathcal{S}_{1}^{c}$.
@@
It is straightforward to see that $\mathcal{T}$ is triangulated. So, we start by showing $\mathcal{T}$ is closed under small coproducts. Let $E_{i}\in\mathcal{T}$ with $i\in I$ for $I$ a set. As $F$ and $G$ both commute with small coproducts, the natural morphism $\alpha\colon\oplus_{i\in I}GF(E_{i})\to GF(\oplus_{i\in I}E_{i})$ is an isomorphism. Also, we have a commutative diagram for each $i\in I$,
$E_{i}$$GF(E_{i})$$\bigoplus_{i\in I}E_{i}$$GF(\bigoplus_{i}E_{i})$$\scriptstyle{\eta_{E_{i}}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{GF(\pi_{i})}$$\scriptstyle{\eta_{\oplus_{i\in I}E_{i}}}$
where $\pi_{i}$ is the natural injection. If $\pi^{\prime}_{i}\colon GF(E_{i})\to\oplus_{i\in I}GF(E_{i})$ is the natural injection, then we know that $GF(\pi_{i})=\alpha\circ\pi_{i}^{\prime}$.
It suffices to show $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Note that $\eta_{\oplus_{i\in I}E_{i}}$ is the unique morphism arising from the universal property of $\oplus_{i\in I}E_{i}$ being a coproduct of the $E_{i}$. Hence, if $GF(\oplus_{i}E_{i})$ with morphisms $GF(\pi_{i})\circ\eta_{E_{i}}$ satisfies the necessary universal property, then $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Consider an $A\in\mathcal{T}$ and morphisms $f_{i}\colon E_{i}\to A$. This gives us morphisms $f_{i}\circ\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$. It can be checked that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\alpha^{-1}\circ GF(\pi_{i})$, and by uniqueness of $f$, we can see that $f\circ\alpha^{-1}$ is the only morphism from $GF(\oplus_{i\in I}E_{i})$ which can satisfy this property. Consequently, we have $GF(\oplus_{i\in I}E_{i})$ satisfying the necessary universal property to show that $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. So, $\mathcal{T}$ is closed under small coproducts.
Now, we check the second claim. If $F$ is fully faithful, then $\eta_{E}$ is an isomorphism for all $E\in\mathcal{S}_{1}$. So, we need to check that $\eta_{E}$ being an isomorphism for every $E\in\mathcal{S}_{1}^{c}$ characterizes fully faithfulness of $F$ on $\mathcal{S}_{1}$. In such a case, we know that $\mathcal{S}_{1}^{c}\subseteq\mathcal{T}$. However, as $\mathcal{T}$ is a triangulated subcategory of $\mathcal{S}_{1}$ that is closed under small coproducts and contains $\mathcal{S}_{1}^{c}$, it follows that $\mathcal{S}_{1}\subseteq\mathcal{T}$. This completes the proof. ∎

### v3

@@ 4. Fully faithfulness & equivalences
Write $\eta$ (resp. $\eta^{\prime}$) for the unit of the adjunction $\Phi_{K}\dashv\Phi_{K^{\prime}}$ (resp. of $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$). By Item˜4.1(2), proving the Proposition amounts to showing the following diagram commutes:
(4.10) $\mathbf{L}t_{1}^{\ast}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\mathbf{L}t_{1}^{\ast}$$\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\mathbf{L}t_{2}^{\ast}\Phi_{K}.$$\scriptstyle{\eta^{\prime}_{\mathbf{L}t_{1}^{\ast}}}$$\scriptstyle{\mathbf{L}t_{1}^{\ast}(\eta)}$$\scriptstyle{\beta^{K^{\prime}}_{\Phi_{K}}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(\beta^{K})}$
As in the proof of Lemma˜4.5, we omit $\mathbf{L}$’s and $\mathbf{R}$’s in the notation for the derived functors to ease notation. Now, the bad news is that Diagram˜4.10 needs not commute. Although, the good news is that it does commute if we replace $\Phi_{(t^{\prime})^{\ast}K}$ by some appropriate isomorphic functor. Specifically, we claim there is an adjunction $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ with unit $\tilde{\eta}^{\prime}$ making Diagram˜4.10 commute (i.e. inscribing $\widetilde{(-)}$ to $\Phi_{(t^{\prime})^{\ast}K}$ and $\eta^{\prime}$) and that moreover there is an isomorphism $\theta:\Phi_{(t^{\prime})^{\ast}K}\cong\widetilde{\Phi}_{(t^{\prime})^{\ast}K}$ satisfying
(4.11) $\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\theta)\circ\eta^{\prime}=\tilde{\eta}^{\prime}$.
Assume that for every $E\in D_{\operatorname{qc}}(Y_{1})$ we have shown:
($\mathcal{P}(E)$) The object $\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$ with the morphism $\Phi_{(t^{\prime})^{\ast}K}(\beta^{K}_{E})\circ\beta^{K^{\prime}}_{\Phi_{K}E}\circ t_{1}^{\ast}(\eta_{E})$ is a reflection of $t_{1}^{\ast}E$ along $\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ (in the sense of [BOR94, Definition 3.1.1]).
Then we can define $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}$ by means of [BOR94, Proposition 3.1.3]. On the one hand, we have $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ with unit $\eta^{\prime}$; thus by [BOR94, Theorem 3.1.5] the pair $(\Phi_{(t^{\prime})^{\ast}K}A,\eta^{\prime}_{A})$ is a reflection of $A\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ along $\Phi_{(t^{\prime})^{\ast}K^{\prime}}$. Now define $\tilde{\eta}^{\prime}=\{\tilde{\eta}^{\prime}_{A}\}_{A\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)}$ by setting $\tilde{\eta}^{\prime}_{A}:=\eta^{\prime}_{A}$ if $A$ is not in the image of $t_{1}^{\ast}:D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ and setting $\tilde{\eta}^{\prime}_{t_{1}^{\ast}E}$, $E\in D_{\operatorname{qc}}(Y_{1})$, to be the unique top morphism in Diagram˜4.10 making it commute. By [BOR94, Proposition 3.1.3], there is a unique functor $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}:D_{\operatorname{qc}}(Y_{1}\times_{S}T)\to D_{\operatorname{qc}}(Y_{2}\times_{S}T)$ whose action on objects coincides with $\Phi_{(t^{\prime})^{\ast}K}$ and such that $\tilde{\eta}^{\prime}:\operatorname{id}\to\widetilde{\Phi}_{(t^{\prime})^{\ast}K^{\prime}}\Phi_{(t^{\prime})^{\ast}K}$ is a natural transformation. Moreover, by [BOR94, Theorem 3.1.5], we have $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ with unit $\tilde{\eta}^{\prime}$. Finally, by [RIE17, Proposition 4.4.1], there is a natural isomorphism $\theta:\Phi_{(t^{\prime})^{\ast}K}\cong\widetilde{\Phi}_{(t^{\prime})^{\ast}K}$ such that (4.11) holds.
It is left for us to show $\mathcal{P}(E)$. So, let $A\in D_{\operatorname{qc}}(Y_{2})$ and consider some $\gamma\colon t_{1}^{\ast}E\to\Phi_{(t^{\prime})^{\ast}K^{\prime}}A$. Our goal is to check that there is a unique morphism $\tilde{\gamma}\colon\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E\to A$ such that the diagram
(4.12) $t_{1}^{\ast}E$$t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}E$$\Phi_{(t^{\prime})^{\ast}K^{\prime}}A$$\Phi_{(t^{\prime})^{\ast}K^{\prime}}t_{2}^{\ast}\Phi_{K}E$$\Phi_{(t^{\prime})^{\ast}K^{\prime}}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$$\scriptstyle{t_{1}^{\ast}(\eta_{E})}$$\scriptstyle{\gamma}$$\scriptstyle{\beta^{K^{\prime}}_{\Phi_{K}(E)}}$$\scriptstyle{\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\beta^{K}_{E})}$$\scriptstyle{\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\tilde{\gamma})}$
commutes. Let $i\in\{1,2\}$. Denote $\zeta^{i}$ to the unit of $t_{i}^{\ast}\dashv(t_{i})_{\ast}$. Given morphisms
$\phi\colon t^{\ast}_{i}M\to B,\quad\psi\colon C\to(t_{i})_{\ast}D,\quad\rho\colon L\to\Phi_{K^{\prime}}P,$
we will use notations
$\phi^{\flat_{i}}\colon M\to(t_{i})_{\ast}B,\quad\psi^{\sharp_{i}}\colon t_{i}^{\ast}C\to D,\quad\rho^{\sharp}\colon\Phi_{K}L\to P$
for the adjunct morphisms of $\phi$, $\psi$, $\rho$ with respect to the adjunctions $t_{i}^{\ast}\dashv(t_{i})_{\ast}$ and $\Phi_{K}\dashv\Phi_{K^{\prime}}$. Suppose there is $\tilde{\gamma}$ turning Diagram˜4.12 into a commutative diagram. Then $\gamma^{\flat_{1}}\colon E\to(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}A$ equals the outer clockwise composite in the following diagram:
$E{{(t_{1})_{\ast}t_{1}^{\ast}E}}$$\Phi_{K^{\prime}}\Phi_{K}E$$(t_{1})_{\ast}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}E$$\Phi_{K^{\prime}}(t_{2})_{\ast}t^{\ast}_{2}\Phi_{K}E$$(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}t_{2}^{\ast}\Phi_{K}E$$\Phi_{K^{\prime}}(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$$(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$$\Phi_{K^{\prime}}(t_{2})_{\ast}A$$(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}A.$$\scriptstyle{\zeta_{E}^{1}}$$\scriptstyle{\eta_{E}}$$\scriptstyle{\delta}$$\scriptstyle{(t_{1})_{\ast}t_{1}^{\ast}(\eta_{E})}$$\scriptstyle{\zeta^{1}_{\Phi_{K^{\prime}}\Phi_{K}E}}$$\scriptstyle{\Phi_{K^{\prime}}(\zeta^{2}_{\Phi_{K}E})}$$\scriptstyle{(t_{1})_{\ast}(\beta^{K^{\prime}}_{\Phi_{K}(E)})}$$\scriptstyle{\alpha^{K^{\prime}}_{t^{\ast}_{2}\Phi_{K}E}}$$\scriptstyle{\Phi_{K^{\prime}}(t_{2})_{\ast}(\beta_{E}^{K})}$$\scriptstyle{(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\beta_{E}^{K})}$$\scriptstyle{\Phi_{K^{\prime}}(t_{2})_{\ast}(\tilde{\gamma})}$$\scriptstyle{(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\tilde{\gamma})}$$\scriptstyle{\alpha_{A}^{K^{\prime}}}$
In this diagram, we note the upper square commutes by naturality of $\zeta^{1}$, the middle square commutes by Lemma˜4.5 (it is Diagram˜4.9 with $K$ replaced by $K^{\prime}$), and the bottom rectangle commutes by naturality of $\alpha^{K^{\prime}}$. Set $\delta$ for the left vertical composite in the last diagram, i.e. $\delta=(\alpha^{K^{\prime}}_{A})^{-1}\circ\gamma^{\flat_{1}}\colon E\to\Phi_{K^{\prime}}(t_{2})_{\ast}A$. Then it follows that
$\delta^{\sharp}\colon\Phi_{K}E\xrightarrow{\zeta^{2}_{\Phi_{K}E}}(t_{2})_{\ast}t_{2}^{\ast}\Phi_{K}E\xrightarrow{(t_{2})_{\ast}(\beta^{K}_{E})}(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E\xrightarrow{(t_{2})_{\ast}(\tilde{\gamma})}(t_{2})_{\ast}A,$
which tells us
$(\delta^{\sharp})^{\sharp_{2}}\colon t_{2}^{\ast}\Phi_{K}E\xrightarrow{\beta_{E}^{K}}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E\xrightarrow{\tilde{\gamma}}A.$
Consequently, we have
$\tilde{\gamma}$ $=(\delta^{\sharp})^{\sharp_{2}}\circ(\beta_{E}^{K})^{-1}$
$=([(\alpha^{K^{\prime}}_{A})^{-1}\circ\gamma^{\flat_{1}}]^{\sharp})^{\sharp_{2}}\circ(\beta_{E}^{K})^{-1},$
which shows uniqueness of $\tilde{\gamma}$. Conversely, defining $\tilde{\gamma}$ by the last formula and reading the previous derivations in reverse order, it follows that Diagram˜4.12 commutes. ∎
Remark 4.8.
Consider the situation as in Proposition 4.7. Denote $\varepsilon$ and $\varepsilon^{\prime}$ respectively for the counits of the adjunctions $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$. From Lemma˜4.1, it follows that (after possibly replacing $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ by an isomorphic functor) the following diagram commutes:
(4.13) $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}$$\mathbf{L}t_{2}^{\ast}\Phi_{K}\Phi_{K^{\prime}}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\mathbf{L}t_{2}^{\ast}$$\mathbf{L}t_{2}^{\ast}.$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\beta^{K^{\prime}})}$$\scriptstyle{\beta^{K}_{\Phi_{K^{\prime}}}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}(\varepsilon)}$$\scriptstyle{\varepsilon^{\prime}_{\mathbf{L}t_{2}^{\ast}}}$
Also, the proof of $\mathcal{P}(E)$ never used that $t$ is affine. Specifically, we showed that with ˜3.1 and given $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ relatively perfect over $Y_{2}$, it is the case that $\mathcal{P}(E)$ holds for $E\in D_{\operatorname{qc}}(Y_{1})$. [^1] The proof of Proposition 4.7 only needs $t$ affine to guarantee that $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$, so that $\mathcal{P}(E)$ plus a replacement of $\Phi_{(t^{\prime})^{\ast}K}$ by a certain isomorphic functor yields a morphism of adjunctions as in the statement of Proposition 4.7. Thus, if it also happens that $t_{1}^{\ast}\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ is essentially surjective, then $\Phi_{(t^{\prime})^{\ast}K^{\prime}}\colon D_{\operatorname{qc}}(Y_{2}\times_{S}T)\to D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ has a left adjoint $L$ (see e.g. [BOR94, Theorem 3.1.5]) such that there is an isomorphism
$L(A)\cong\Phi_{(t^{\prime})^{\ast}K}(A)$
for each $A\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$. However, one should be careful, for the isomorphism above might not be natural in $A$.
Proposition 4.9.
Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ where $t$ is one of the following:
  1. (1)
    $t$ is affine and faithfully flat (e.g. $t=1_{S}$)
  2. (2)
    $t\colon\operatorname{Spec}(k)\to S$ is any morphism from a field (so here, we want the condition to all for all such morphisms)
  3. (3)
    $t\colon\operatorname{Spec}(\kappa(p))\to S$ is any closed immersion associated to a closed point $p\in S$ (so here, we want the condition to all for all such morphisms),
then so is $\Phi_{K}$.
@@
Set $K^{\prime}$ to be the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. By Theorem˜3.6, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$ for each $t\colon T\to S$ in the statement. Hence, Proposition˜3.8 tells us that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ for each $t$ in the statement. So, from Corollary˜3.9, we know that $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D^{b}_{\operatorname{coh}}$ where $t$ is any morphism in the statement.
We start with the first case. Denote $\eta$ and $\varepsilon$ (resp. $\eta^{\prime}$, $\varepsilon^{\prime}$) for the unit and counit of the adjunction $\Phi_{K}\dashv\Phi_{K^{\prime}}$ (resp. of $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$). Since $t$ is faithfully flat, so are each $t_{i}$; hence each $\mathbf{L}(t_{i})^{\ast}$ reflects isomorphisms. Therefore, by commutativity of Diagram˜4.10 (resp. by Diagram˜4.13), it follows that $\eta^{\prime}$ (resp. $\varepsilon^{\prime}$) being an isomorphism implies $\eta$ (resp. $\varepsilon$) is an isomorphism. The result then follows from [STA25, Tag 07RB].
Now, we check the second and third case. In both cases, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ with $t\colon\operatorname{Spec}(\kappa(p))\to S$ is any closed immersion associated to a closed point $p\in S$. In this case, one can argue exactly as in [HLS09, Proposition 2.15]. Indeed, we only need to replace Proposition 2.9 of loc. cit. with Proposition˜3.8, i.e. ‘locally projective’ constraints of loc. cit. can be replaced with ‘proper’ in light of our work. ∎